You are buying flowers to hand out at a school dance. Roses cost $30 for a dozen but cost more if bought individually. With the money you have, you can buy 7 dozen and 4 single roses, or 64 single roses. How much is one rose? How much money do you have?
Answer by hydromojo(20)

Let the price of single roses be $X
With the money you have, you can buy
1) 7 dozen roses + 4 single roses, OR
2) 64 single roses
Since the same money can buy the roses, we can equate both 1 & 2, so
7 dozen roses + 4 singles = 64 single roses
i.e. 7 dozen roses = (64-4) singles
i.e. 7 dozens = 60 singles
now, we know the price of a dozen ($30), so 7 dozens = $210
so, 60 singles also cost $210, thus each single costs 210/60 = $3.5
So, you have money for 7dozen roses + 4 singles
= $210 + 4 * 3.5 = $224
and each single costs $3.5